# L4DR: LiDAR-4DRadar Fusion for Weather-Robust 3D Object Detection

### v1

@@ Abstract
LiDAR-based vision systems are integral for 3D object detection, which is crucial for autonomous navigation. However, they suffer from performance degradation in adverse weather conditions due to the quality deterioration of LiDAR point clouds. Fusing LiDAR with the weather-robust 4D radar sensor is expected to solve this problem. However, the fusion of LiDAR and 4D radar is challenging because they differ significantly in terms of data quality and the degree of degradation in adverse weather. To address these issues, we introduce L4DR, a weather-robust 3D object detection method that effectively achieves LiDAR and 4D Radar fusion. Our L4DR includes Multi-Modal Encoding (MME) and Foreground-Aware Denoising (FAD) technique to reconcile sensor gaps, which is the first exploration of the complementarity of early fusion between LiDAR and 4D radar. Additionally, we design an Inter-Modal and Intra-Modal ({IM}2) parallel feature extraction backbone coupled with a Multi-Scale Gated Fusion (MSGF) module to counteract the varying degrees of sensor degradation under adverse weather conditions. Experimental evaluation on a VoD dataset with simulated fog proves that L4DR is more adaptable to changing weather conditions. It delivers a significant performance increase under different fog levels, improving the 3D mAP by up to 19.99% over the traditional LiDAR-only approach. Moreover, the results on the K-Radar dataset validate the consistent performance improvement of L4DR in real-world adverse weather conditions.
@@ Introduction
3D object detection is fundamental to vision systems of unmanned platforms, extensively utilized in applications such as intelligent robots navigation (Ghita et al. 2024; Xu et al. 2022) and autonomous driving (Bijelic et al. 2020). Full Driving Automation (FDA, Level 5) relies on weather-robust 3D object detection, providing precise 3D bounding boxes even under various challenging adverse weather conditions (Qian et al. 2021). Owing to the high resolution and strong interference resistance of LiDAR sensors, LiDAR-based 3D object detection has emerged as a mainstream area of research (Yan et al. 2023; Xia et al. 2023; Wu et al. 2024). However, LiDAR sensors exhibit considerable sensitivity to weather conditions. In adverse scenarios such as rain and fog, the scanning signals suffer from substantial degradation and increased noise (Huang et al. 2024; Hahner et al. 2021). This degradation can negatively impact 3D detectors, compromising the reliability of autonomous perception systems.
@@
Aside from LiDAR, 4D (range, azimuth, Doppler, and elevation) millimeter-wave radar has increasingly been recognized (Han et al. 2023; Sun and Zhang 2021). As shown in Figure 1 (a), 4D radar outperforms LiDAR in weather robustness, velocity measurement, and detection range. The millimeter-wave signals of 4D radar have wavelengths much larger than the tiny particles that form fog, rain, and snow (Golovachev et al. 2018; Balal, Pinhasi, and Pinhasi 2016), exhibiting reduced susceptibility to weather disturbances. As shown in Fig. 1 (b), the performance gap between LiDAR and 4D radar decreases as the severity of the weather rises. Therefore, the 4D radar sensor is suitable for various weather conditions. However, LiDAR is still far ahead of the radar in the important metrics of object classification and resolution. These circumstances make it very feasible to promote the full fusion of 4D radar and LiDAR data to improve 3D object detection. Pioneering approaches such as InterFusion (Wang et al. 2022b), M2Fusion (Wang et al. 2022a), and 3D-LRF (Chae, Kim, and Yoon 2024) represent the initial explorations into the fusion of LiDAR and 4D radar, showing significant performance improvements compared to models employing LiDAR-only or 4D radar-only.
Figure 2: The challenges of fusing LiDAR and 4D radar in adverse weather conditions. (a) Significant quality disparity between the LiDAR and the 4D radar. (b) Severe degradation of LiDAR data quality in adverse weather.
Despite these advances, the fusion of LiDAR and 4D radar sensors has not addressed the substantial disparities in data quality and the different degrees of degradation in adverse weather. As shown in Figure 2 (a), a primary challenge arises from the significant quality disparity between the LiDAR sensor and the 4D radar sensor. A second challenge pertains to varying sensor degradation under adverse weather conditions. Figure 2 (b) shows LiDAR sensors undergo severe data degradation in adverse weather. Conversely, the data quality decrease in 4D radar is significantly lower (Han et al. 2023; Sun and Zhang 2021) (Details of real-world LiDAR and 4D radar point cloud distributions under different weather conditions are given in Supplementary Material A.1.). This motivates the consideration of emphasizing LiDAR and 4D radar sensors with varying degrees while considering weather conditions during data fusion. The substantial data quality disparities and different degrees of sensor degradation is not properly addressed in existing LiDAR and 4D radar fusion methods in adverse weather conditions.
@@
To address the challenges above, we propose L4DR, a weather-robust 3D object detection via LiDAR and 4D radar fusion. We designed a Multi-Modal Encoder (MME) module and a Foreground-Aware Denoising (FAD) to address the substantial data quality disparities. This innovation pioneers the use of early data-level fusion strategies to mitigate the disparities between LiDAR and 4D radar while concurrently improving the data quality from each modality. In response to the varying degrees of sensor degradation under diverse weather conditions, L4DR incorporates the Inter-Modal and Intra-Modal ({IM}2) backbone and the Multi-Scale Gated Fusion (MSGF) module, adaptively focusing on the significant modal features in different weather conditions.
In Figure  3, our comprehensive testing on VoD (Palffy et al. 2022) and K-Radar datasets (Paek, Kong, and Wijaya 2022) showcase the resilience and superior performance of L4DR across various simulated and real-world adverse weather disturbances. Our main contributions are as follows:
  • We introduce the innovative Multi-Modal Encoder (MME) module, which achieves LiDAR and 4D radar fusion without resorting to error-prone processes (e.g., depth estimation, height compression). This approach effectively bridges the substantial LiDAR and 4D radar data quality disparities.
  • We designed an {IM}2 backbone with a Multi-Scale Gated Fusion (MSGF) module, adaptively extracting salient features from LiDAR and 4D radar sensors under various weather conditions. This enables the model to adapt to varying levels of sensor degradation under adverse weather conditions.
@@
Researchers have developed single-stage and two-stage methods to tackle challenges for 3D object detection. Single-stage detectors such as VoxelNet (Yan, Mao, and Li 2018), PointPillars (Lang et al. 2019), 3DSSD (Yang et al. 2020), DSVT (Wang et al. 2023a) utilize PointNet++ (Qi et al. 2017), sparse convolution, or other point feature encoder to extract features from point clouds and perform detection in the Bird’s Eye View (BEV) space. Conversely, methods such as PV-RCNN (Shi et al. 2020), PV-RCNN++ (Shi et al. 2022), Voxel-RCNN (Deng et al. 2021), and VirConv  (Wu et al. 2023) focus on two-stage object detection, integrating RCNN networks into 3D detectors. Even though these mainstream methods have gained excellent performance in normal weather, they lack robustness and still need improvement under various adverse weather conditions.
@@
LiDAR sensors may undergo degradation under adverse weather conditions such as snow, fog, and rain. Physics-based simulations (Teufel et al. 2022; Hahner et al. 2022, 2021; Kilic et al. 2021) have been explored to reproduce point clouds under adverse weather to alleviate the issue of data scarcity. (Charron, Phillips, and Waslander 2018; Heinzler et al. 2020) utilized the DROR, DSOR, or convolutional neural networks (CNNs) to classify and filter LiDAR noise points. (Xu et al. 2021) designed a general completion framework that addresses the problem of domain adaptation across different weather conditions. (Huang et al. 2024) designed a general knowledge distillation framework that transfers sunny performance to rainy performance. However, these methods primarily rely on single-LiDAR modal data, which will be constrained by the decline in the quality of LiDAR under adverse weather conditions.
@@
LiDAR-radar fusion for 3D object detection has gained increasing attention in recent years. MVDNet (Qian et al. 2021) has designed a framework for fusing LiDAR and radar. ST-MVDNet (Li et al. 2022) and ST-MVDNet++ (Li, O’Toole, and Kitani 2023) incorporate a self-training teacher-student to MVDNet to enhance the model. Bi-LRFusion (Wang et al. 2023b) framework employs a bidirectional fusion strategy to improve dynamic object detection performance. However, these studies only focus on 3D radar and LiDAR. As research progresses, the newest studies continue to drive the development of LiDAR-4D radar fusion. M2-Fusion (Wang et al. 2022a) and InterFusion (Wang et al. 2022b) explore new LiDAR and 4D radar fusion approaches. However, these methods have not considered and overcome the challenges of fusing 4D radar and LiDAR under adverse weather conditions to promote weather-robust 3D object detection.
@@
For an outdoor scene, we denote a LiDAR point cloud as $\mathcal{P}^{l}=\{p^{l}_{i}\}_{i=1}^{N_{l}}$, $p^{l}_{i}=\{x_{i},y_{i},z_{i},\{C^{l}_{i}\}\}$ and a 4D radar point cloud as $\mathcal{P}^{r}=\{p^{r}_{i}\}_{i=1}^{N_{r}}$, $p^{r}_{i}=\{x_{i},y_{i},z_{i},\{C^{r}_{i}\}\}$, where $i$ indexes individual points, $x,y,z$ denote 3D world coordinates, and $N^{m}$ and $C^{m}$ denote the number of points and the point-wise initial data features of modality $m\in\{l,r\}$ ($l$ for LiDAR and $r$ for radar). Subsequently, a multi-modal model $\mathcal{M}$ extract deep features $\mathcal{F}^{m}$ from $\mathcal{P}^{m}$, written as $\mathcal{F}^{m}=g(f_{\mathcal{M}}(\mathcal{P}^{m};\Theta))$. The objective of 3D object detection is to regress the 3D bounding boxes $B=\{b_{i}\}_{i=1}^{N_{b}}$, $B\in\mathbb{R}^{N_{b}\times 7}$, using fused feature $\mathcal{F}^{f}=\phi(\mathcal{F}^{l},\mathcal{F}^{r})$, where $\phi$ donates fusion method. Each box $b_{i}$ is encoded using 3D coordinates, size, and orientation, denoted by $b_{i}=\{x_{i},y_{i},z_{i},l_{i},w_{i},h_{i},\theta_{i}\}$.
Significant quality disparities between LiDAR and 4D radar sensors
. As mentioned before, there is a huge difference between $\mathcal{P}^{l}$ and $\mathcal{P}^{r}$ in the same scene. Specifically, we use the Chamfer Distance ($d_{CD}$) (Borgefors 1986) to quantitatively measure the point cloud distance between $\mathcal{P}^{l}$ and $\mathcal{P}^{r}$ in the VoD (Palffy et al. 2022) dataset.
Figure 4: L4DR framework. (a) Multi-Modal Encoder (MME) fuses bi-directional data for both LiDAR and 4D radar modalities at the Encoder stage to obtain higher quality BEV features. Next, (b) Foreground-Aware Denoising (FAD) performs denoising by segmenting foreground semantics per 4D radar point. Finally, (c) Inter-Modal and Intra-Modal ({IM}2) backbone coupled with Multi-Scale Gated Fusion (MSGF) uses a gating strategy to filter features to avoid redundant information while extracting inter-modal and intra-modal features in parallel.
The result of $d_{CD}(\mathcal{P}^{l},\mathcal{P}^{r})$ is 342.6 m between the LiDAR and 4D Radar point clouds even in the same scene, indicating a significant difference between $\mathcal{P}^{l}$ and $\mathcal{P}^{r}$. This data quality disparities phenomenon poses a significant challenge to data fusion.
@@
To fully fuse two modalities, one possible solution is to use $P^{l}$ to enhance the highly sparse $P^{r}$ that lacks discriminative details. Therefore, our L4DR includes a Multi-Modal Encoder (MME, Figure 4(a)), which performs early-fusion complementarity at the data encoder level. However, we found that direct data fusion would also cause much of the noise in $\mathcal{P}^{r}$ to spread to $\mathcal{P}^{l}$. Therefore, we integrated Foreground-Aware Denoising (FAD, Figure 4(b)) into L4DR before MME to filter out most of the noise in $\mathcal{P}^{r}$.
Significant degradation of LiDAR sensor in adverse weather conditions
. Compared to 4D radar, the quality of LiDAR point cloud $P^{l}$ is more easily affected by adverse weather conditions, leading to lower data quality and varying feature presentations $\mathcal{F}^{l}$. Taking foggy weather as an example, we conducted four levels of fog simulations (Hahner et al. 2021) (donated as $\psi$) on the VoD dataset to simulate different degrees of foggy weather. The simulated LiDAR data is denoted as $P^{l}_{\mathcal{L}}=\psi(P^{l},\mathcal{L})$, $\mathcal{L}\in[0,4]$ indicating the level of fog simulation with zero representing clear weather. Subsequently, cosine similarity $\mathcal{S}_{cos}$ was used to quantitatively measure the similarity between features $\mathcal{F}^{l}_{\mathcal{L}}$ under different levels of fog simulation $\mathcal{L}$, to the features at fog level zero $\mathcal{F}^{l}_{0}$ in the same scene. The cosine similarity is given as
$\mathcal{S}_{cos}(\mathcal{F}^{l}_{0},\mathcal{F}^{l}_{\mathcal{L}})=\frac{\mathcal{F}^{l}_{0}\cdot\mathcal{F}^{l}_{\mathcal{L}}}{||\mathcal{F}^{l}_{0}||\times||\mathcal{F}^{l}_{\mathcal{L}}||},$ (1)
The results in Table 1 show that the degree of variation in $\mathcal{F}^{l}$ is significantly affected by fog level $\mathcal{L}$ in different convolutional layers $\mathcal{D}$. In real-world scenarios where weather conditions are subject to constant change, the feature presentation $\mathcal{F}^{l}$ is frequently varying, which will also affect the presentation of the fusion feature $\mathcal{F}^{f}$. This phenomenon has a detrimental impact on the weather-robust stability advantage of $\mathcal{F}^{r}$, which in turn impedes the realization of a weather-robust 3D detector.
Layer Index $\overline{\mathcal{S}_{cos}}(\mathcal{F}^{l}_{0},\mathcal{F}^{l}_{1})$ $\overline{\mathcal{S}_{cos}}(\mathcal{F}^{l}_{0},\mathcal{F}^{l}_{2})$ $\overline{\mathcal{S}_{cos}}(\mathcal{F}^{l}_{0},\mathcal{F}^{l}_{3}))$ $\overline{\mathcal{S}_{cos}}(\mathcal{F}^{l}_{0},\mathcal{F}^{l}_{4})$
$\mathcal{D}=0$ (BEV) 0.84($\downarrow$0.16) 0.64($\downarrow$0.36) 0.52($\downarrow$0.48) 0.27($\downarrow$0.73)
$\mathcal{D}=1$ 0.95($\downarrow$0.05) 0.83($\downarrow$0.17) 0.72($\downarrow$0.28) 0.57($\downarrow$0.43)
$\mathcal{D}=2$ 0.98($\downarrow$0.02) 0.87($\downarrow$0.13) 0.77($\downarrow$0.23) 0.73($\downarrow$0.27)
$\mathcal{D}=3$ 0.99($\downarrow$0.01) 0.93($\downarrow$0.07) 0.71($\downarrow$0.29) 0.65($\downarrow$0.35)
Table 1: The cosine similarity of VoD dataset LiDAR features under different concentrations of adverse weather with respect to features under clear weather.
@@
Previous backbones focusing solely on fused inter-modal features $\mathcal{F}^{f}$ overlook the strengths of each modality, leading to challenges in addressing the frequent fluctuation of $\mathcal{F}^{l}$. To address this issue and ensure robust fusion across diverse weather conditions, we introduce the Inter-Modal and Intra-Modal ({IM}2, Figure 4 (c)) backbone. This design simultaneously emphasizes inter-modal and intra-modal features, enhancing model adaptability. However, redundancy between these features arises. Drawing inspiration from gated fusion techniques (Hosseinpour, Samadzadegan, and Javan 2022; Song, Zhao, and Skinner 2024a), we propose the Multi-Scale Gated Fusion (MSGF, Figure 4 (d)) module. MSGF utilizes inter-modal features $\mathcal{F}^{f}$ to filter intra-modal features $\mathcal{F}^{l}$ and $\mathcal{F}^{r}$, effectively reducing feature redundancy. Our experimental results in Table. 5 shows the {IM}2 and MSGF modules significantly enhance the model’s capability to adapt to varying adverse weather conditions, ensuring the ability of weather-robustness.
@@
Due to multipath effects, 4D radar contains significant noise points. Despite applying the Constant False Alarm Rate (CFAR) algorithm to filter out noise during the data acquisition process, the noise level remains substantial. It is imperative to further reduce the clutter noise in 4D radar data before early data fusion to avoid noise spreading. Considering the minimal contribution of background points to object detection, this work introduces point-level foreground semantic segmentation to 4D radar denoising, performing a Foreground-Aware Denoising. Specifically, we first utilize PointNet++ (Qi et al. 2017) combined with a segmentation head as $\chi$, to predict the foreground semantic probability $\mathcal{S}=\chi(\mathcal{P}^{r})$ for each point in the 4D radar. Subsequently, points with a foreground probability below a predefined threshold $\tau$ are filtered out, that is $\mathcal{P}^{r}_{new}=\{p^{r}_{i}|\mathcal{S}_{i}\geq\tau\}$. FAD effectively filters out as many noise points as possible while preventing the loss of foreground points.
@@
As illustrated in Figure 5, we first refer (Lang et al. 2019) to encode the LiDAR point cloud into a pillar set $P^{l}=\{p^{l}_{i}\}_{i=1}^{N}$. Each LiDAR point $p^{l}_{(i,j)}$ in pillar $p^{l}_{i}$ encoded using encoding feature $\boldsymbol{f}^{l}_{(i,j)}$ as
@@
The feature propagation is beneficial because $\lambda$ and $\Omega$ are helpful for object classification, while Doppler information $\boldsymbol{\mathcal{V}}$ is crucial for distinguishing dynamic objects (Song, Zhao, and Skinner 2024b). Cross-modal feature sharing makes comprehensive use of these advantages and cross-modal offsets [$\boldsymbol{\mathcal{Y}}^{m}_{cl},\boldsymbol{\mathcal{Y}}^{m}_{cr}$], $m\in\{l,r\}$ also further enrich the geometric information. This MME method compensates for the data quality of 4D radar under normal weather conditions and can also enhance the quality of LiDAR in adverse weather, providing richer information for feature extraction. Subsequently, we applied a linear layer and max pooling operations to the fused pillar encoding features $\widehat{\boldsymbol{f}}$ to obtain the corresponding modal BEV features $\mathcal{F}$.

### v3

@@ Abstract
LiDAR-based 3D object detection is crucial for autonomous driving. However, due to the quality deterioration of LiDAR point clouds, it suffers from performance degradation in adverse weather conditions. Fusing LiDAR with the weather-robust 4D radar sensor is expected to solve this problem; however, it faces challenges of significant differences in terms of data quality and the degree of degradation in adverse weather. To address these issues, we introduce L4DR, a weather-robust 3D object detection method that effectively achieves LiDAR and 4D Radar fusion. Our L4DR proposes Multi-Modal Encoding (MME) and Foreground-Aware Denoising (FAD) modules to reconcile sensor gaps, which is the first exploration of the complementarity of early fusion between LiDAR and 4D radar. Additionally, we design an Inter-Modal and Intra-Modal ({IM}2) parallel feature extraction backbone coupled with a Multi-Scale Gated Fusion (MSGF) module to counteract the varying degrees of sensor degradation under adverse weather conditions. Experimental evaluation on a VoD dataset with simulated fog proves that L4DR is more adaptable to changing weather conditions. It delivers a significant performance increase under different fog levels, improving the 3D mAP by up to 20.0% over the traditional LiDAR-only approach. Moreover, the results on the K-Radar dataset validate the consistent performance improvement of L4DR in real-world adverse weather conditions.
@@ Introduction
3D object detection is a fundamental vision task of unmanned platforms, extensively utilized in applications such as intelligent robot navigation (Ghita et al. 2024; Xu et al. 2022) and autonomous driving (Bijelic et al. 2020). For example, Full Driving Automation (FDA, Level 5) relies on weather-robust 3D object detection, which provides precise 3D bounding boxes even under various challenging adverse weather conditions (Qian et al. 2021). Owing to the high resolution and strong interference resistance of LiDAR sensors, LiDAR-based 3D object detection has emerged as a mainstream area of research (Yan et al. 2023; Xia et al. 2023; Wu et al. 2024). However, LiDAR sensors exhibit considerable sensitivity to weather conditions. In adverse scenarios, the scanned LiDAR point clouds suffer from substantial degradation and increased noise (Huang et al. 2024; Hahner et al. 2021). This degradation negatively impacts 3D detectors, compromising the reliability of autonomous perception systems.
Aside from LiDAR, 4D (range, azimuth, Doppler, and elevation) millimeter-wave radar is gaining recognition as a vital perception sensor (Han et al. 2023; Sun and Zhang 2021). As shown in Fig.1 (a), 4D radar outperforms LiDAR in weather robustness, velocity measurement, and detection range. The millimeter-wave signals of 4D radar have wavelengths much larger than the tiny particles in fog, rain, and snow (Golovachev et al. 2018; Balal, Pinhasi, and Pinhasi 2016), exhibiting reduced susceptibility to weather disturbances. As shown in Fig.1 (b), the performance gap between LiDAR and 4D radar decreases as the severity of the weather rises. Hence, the 4D radar sensor is more suitable for adverse weather conditions than the LiDAR sensor. However, LiDAR has advantages in terms of higher distance and angular point resolution. These circumstances make it feasible to promote the full fusion of 4D radar and LiDAR data to improve 3D object detection. Pioneering approaches such as InterFusion (Wang et al. 2022b), M2Fusion (Wang et al. 2022a), and 3D-LRF (Chae, Kim, and Yoon 2024) conducted initial attempts to fuse LiDAR and 4D radar data, have demonstrated significant performance improvements over single-sensor models.
@@
Figure 2: (a) Significant quality disparity between the LiDAR and the 4D radar. (b) Severe degradation of LiDAR data quality in adverse weather. (c) Comparison of previous LiDAR-4DRadar fusion and our fusion, highlighting our innovative framework designs to address challenges (a) and (b).
Despite progress, the significant data quality disparities and varying sensor degradation in adverse weather conditions remain largely unaddressed. As shown in Fig.2 (a), a primary challenge arises from the significant quality disparity between the LiDAR and 4D radar sensors. A second challenge pertains to varying sensor degradation under adverse weather conditions. Fig.2 (b) shows LiDAR sensors undergo severe data degradation in adverse weather. Conversely, the data quality decrease in 4D radar is significantly lower (Han et al. 2023; Sun and Zhang 2021) (Details are given in supplementary material). The varying degradation of sensors leads to fluctuations in the fused features, resulting in difficulty for the model to maintain high performance under adverse weather conditions. Therefore, developing a weather-robust 3D object detector requires overcoming the challenges of significant quality disparities and varying sensor degradation.
To address the above challenges, we propose L4DR, a novel two-stage fusion framework for LiDAR and 4D radar (shown in Fig. 2 (c)). The first-stage fusion is performed by the 3D Data Fusion Encoder, which contains the Foreground-Aware Denoising (FAD) module and the Multi-Modal Encoder (MME) module. This first-stage data fusion augments LiDAR and 4D radar with each other to tackle the challenge of significant data quality disparities between LiDAR and 4D radar. The second-stage Feature Fusion is to fuse features between 2D Backbones to address the second challenge of the varying sensor degradation. It consists of the Inter-Modal and Intra-Modal ({IM}2) backbone and the Multi-Scale Gated Fusion (MSGF) module. Different from traditional methods that fuse features before extraction on the left side of Fig.2 (c), L4DR integrates a continuous fusion process throughout the feature extraction, adaptively focusing on the significant modal features in different weather conditions.
@@
In Fig. 3, our comprehensive experimental results showcase the superior performance of L4DR across various simulated and real-world adverse weather disturbances. Our main contributions are as follows:
  • We introduce the innovative Multi-Modal Encoder (MME) module, which achieves LiDAR and 4D radar data early fusion without resorting to error-prone processes (e.g., depth estimation). It effectively bridges the substantial LiDAR and 4D radar data quality disparities.
  • We designed an {IM}2 backbone with a Multi-Scale Gated Fusion (MSGF) module, adaptively extracting salient features from LiDAR and 4D radar in different weather conditions. This enables the model to adapt to varying levels of sensor degradation under adverse weather conditions.
@@
Figure 4: L4DR framework. (a) Foreground-Aware Denoising (FAD) performs denoising by segmenting foreground semantics per 4D radar point. Next, (b) Multi-Modal Encoder (MME) fuses bi-directional data for both LiDAR and 4D radar modalities at the Encoder stage to obtain higher quality BEV features. Finally, (c) Inter-Modal and Intra-Modal ({IM}2) backbone coupled with Multi-Scale Gated Fusion (MSGF) uses a gating strategy to filter features to avoid redundant information while extracting inter-modal and intra-modal features in parallel.
@@
Researchers have developed single-stage and two-stage methods to tackle challenges for 3D object detection. Single-stage detectors such as VoxelNet (Yan, Mao, and Li 2018), PointPillars (Lang et al. 2019), 3DSSD (Yang et al. 2020), DSVT (Wang et al. 2023a) utilize PointNet++ (Qi et al. 2017), sparse convolution, or other point feature encoder to extract features from point clouds and perform detection in the Bird’s Eye View (BEV) space. Conversely, methods such as PV-RCNN (Shi et al. 2020), PV-RCNN++ (Shi et al. 2022), Voxel-RCNN (Deng et al. 2021), and VirConv  (Wu et al. 2023) focus on two-stage object detection, integrating RCNN networks into 3D detectors. Even though these mainstream methods have gained excellent performance in normal weather, they still lack robustness under various adverse weather conditions.
@@
LiDAR sensors may undergo degradation under adverse weather conditions. Physics-based simulations (Teufel et al. 2022; Hahner et al. 2022, 2021; Kilic et al. 2021) have been explored to reproduce point clouds under adverse weather to alleviate the issue of data scarcity. (Charron, Phillips, and Waslander 2018; Heinzler et al. 2020) utilized the DROR, DSOR, or convolutional neural networks (CNNs) to classify and filter LiDAR noise points. (Xu et al. 2021) designed a general completion framework that addresses the problem of domain adaptation across different weather conditions. (Huang et al. 2024) designed a general knowledge distillation framework that transfers sunny performance to rainy performance. However, these methods primarily rely on single-LiDAR modal data, which will be constrained by the decline in the quality of LiDAR under adverse weather conditions.
@@
MVDNet (Qian et al. 2021) has designed a framework for fusing LiDAR and radar. ST-MVDNet (Li et al. 2022) and ST-MVDNet++ (Li, O’Toole, and Kitani 2023) incorporate a self-training teacher-student to MVDNet to enhance the model. Bi-LRFusion (Wang et al. 2023b) framework employs a bidirectional fusion strategy to improve dynamic object detection performance. However, these studies only focus on 3D radar and LiDAR. As research progresses, the newest studies continue to drive the development of LiDAR-4D radar fusion. M2-Fusion (Wang et al. 2022a), InterFusion (Wang et al. 2022b), and 3D-LRF (Chae, Kim, and Yoon 2024) explore LiDAR and 4D radar fusion. However, these methods have not considered and overcome the challenges of fusing 4D radar and LiDAR under adverse weather conditions.
@@
For an outdoor scene, we denote LiDAR point cloud as $\mathcal{P}^{l}=\{p^{l}_{i}\}_{i=1}^{N_{l}}$ and 4D radar point cloud as $\mathcal{P}^{r}=\{p^{r}_{i}\}_{i=1}^{N_{r}}$, where $p$ denotes 3D points. Subsequently, a multi-modal model $\mathcal{M}$ extract deep features $\mathcal{F}^{m}$ from $\mathcal{P}^{m}$, written as $\mathcal{F}^{m}=g(f_{\mathcal{M}}(\mathcal{P}^{m};\Theta))$, where $m\in\{l,r\}$. The fusion features are then obtained by $\mathcal{F}^{f}=\phi(\mathcal{F}^{l},\mathcal{F}^{r})$, where $\phi$ donates fusion method. The objective of 3D object detection is to regress the 3D bounding boxes $B=\{b_{i}\}_{i=1}^{N_{b}}$, $B\in\mathbb{R}^{N_{b}\times 7}$.
Significant data quality disparity.
@@
As mentioned before, there is a dramatic difference between $\mathcal{P}^{l}$ and $\mathcal{P}^{r}$ in the same scene. To fully fuse two modalities, we can use $P^{l}$ to enhance the highly sparse $P^{r}$ that lacks discriminative details. Therefore, our L4DR includes a Multi-Modal Encoder (MME, Figure 4 (b)), which performs data early-fusion complementarity at the encoder. However, we found that direct data fusion would also cause noises in $\mathcal{P}^{r}$ to spread to $\mathcal{P}^{l}$. Therefore, we integrated Foreground-Aware Denoising (FAD, Figure 4 (a)) into L4DR before MME to filter out most of the noise in $\mathcal{P}^{r}$.
Varying sensor degradation in adverse weather.
@@
Compared to 4D radar, the quality of LiDAR point cloud $P^{l}$ is more easily affected by adverse weather conditions, leading to varying feature presentations $\mathcal{F}^{l}$. Previous backbones focusing solely on fused inter-modal features $\mathcal{F}^{f}$ overlook the weather robustness of 4D radar, leading to challenges in addressing the frequent fluctuation of $\mathcal{F}^{l}$. To ensure robust fusion across diverse weather conditions, we introduce the Inter-Modal and Intra-Modal ({IM}2, Figure 4 (c)) backbone. This design simultaneously focuses on inter-modal and intra-modal features, enhancing model adaptability. However, redundancy between these features arises. Inspired by gated fusion techniques (Hosseinpour, Samadzadegan, and Javan 2022; Song, Zhao, and Skinner 2024a), we propose the Multi-Scale Gated Fusion (MSGF, Figure 4 (d)) module. MSGF utilizes inter-modal features $\mathcal{F}^{f}$ to filter intra-modal features $\mathcal{F}^{l}$ and $\mathcal{F}^{r}$, effectively reducing feature redundancy.
@@
Due to multipath effects, 4D radar contains significant noise points. Despite applying the Constant False Alarm Rate (CFAR) algorithm to filter out noise during the data acquisition process, the noise level remains substantial. It is imperative to further reduce clutter noise in 4D radar data before early data fusion to avoid spreading noise. Considering the minimal contribution of background points to object detection, this work introduces point-level foreground semantic segmentation to 4D radar denoising, performing a Foreground-Aware Denoising. Specifically, we first utilize PointNet++ (Qi et al. 2017) combined with a segmentation head as $\chi$, to predict the foreground semantic probability $\mathcal{S}=\chi(\mathcal{P}^{r})$ for each point in the 4D radar. Subsequently, points with a foreground probability below a predefined threshold $\tau$ are filtered out, that is $\mathcal{P}^{r}_{new}=\{p^{r}_{i}|\mathcal{S}_{i}\geq\tau\}$. FAD effectively filters out as many noise points as possible while preventing the loss of foreground points.
@@
As illustrated in Figure 5, we innovate the traditional unimodal Pillars coding into multimodal Pillars coding to perform initial fusion at the data level, extracting richer information for subsequent feature processing. Firstly, referring (Lang et al. 2019), we encode the LiDAR point cloud into a pillar set $P^{l}=\{p^{l}_{i}\}_{i=1}^{N}$. Each LiDAR point $p^{l}_{(i,j)}$ in pillar $p^{l}_{i}$ encoded using encoding feature $\boldsymbol{f}^{l}_{(i,j)}$ as
@@
The feature propagation is beneficial because $\lambda$ and $\Omega$ are helpful for object classification, while Doppler information $\boldsymbol{\mathcal{V}}$ is crucial for distinguishing dynamic objects (Song, Zhao, and Skinner 2024b). Cross-modal feature sharing makes comprehensive use of these advantages and cross-modal offsets [$\boldsymbol{\mathcal{Y}}^{m}_{cl},\boldsymbol{\mathcal{Y}}^{m}_{cr}$], $m\in\{l,r\}$ also further enrich the geometric information. This MME method compensates for the data quality of 4D radar under normal weather conditions and can also enhance the quality of LiDAR in adverse weather. Subsequently, we applied a linear layer and max pooling operations to the fused pillar encoding features $\widehat{\boldsymbol{f}}$ to obtain the corresponding modal BEV features $\mathcal{F}$.
@@
Loss Function and Training Strategy.
We trained our L4DR with the following losses :
$\mathcal{L}_{cls}+\beta_{loc}\mathcal{L}_{loc}+\beta_{fad}\mathcal{L}_{fad}),$ (7)
where $\beta_{cls}$ = 1, $\beta_{loc}$ = 2, $\beta_{cls}$ = 0.5, the $\mathcal{L}_{fad}$ is object classification focal loss, the $\mathcal{L}_{loc}$ is object localization regression loss, and the $\mathcal{L}_{fad}$ is focal classification loss of 4D radar foreground point in FAD module. We use Adam optimizer with lr = 1e-3, $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.999.